OpenStudy (anonymous):

Find the slope of the line that contains the points (-5, 3) and (-1, -5). A: -4 B:-2 C: 2 D: 4

OpenStudy (rizags):

do you know the slope equation?

OpenStudy (anonymous):

(-5-3)/(-1-(-5))

OpenStudy (anonymous):

No :( @Rizags

Mehek (mehek14):

\(\dfrac{y_2 - y_1}{x_2 - x_1}\) plug in both ordered pairs \(x_1\) \(y_1\) \(x_2\) \(y_2\) (-5, 3) (-1, -5) \(\dfrac{-5 - 3}{-1 - (-5)} = \dfrac{-8}{4} = -2\)

OpenStudy (anonymous):

thank you so much! @Mehek14

Mehek (mehek14):

yw \(\ddot\smile\)